OpenStudy (anonymous):

Write a quadratic variation equation if g(x) varies directly with x^2 and g(x) = 144 when x = 6. A. G(x)=2x^2 B. G(x) = 6x^2 C. G(x)= 4x^2 D. G(x) = 1/4x ^2

random231 (random231):

g(x) is proportional to x^2 so g(x)=k*x^2 can you find k?

OpenStudy (dariusx):

c

OpenStudy (anonymous):

I don't really know how to do any of it :(

OpenStudy (dariusx):

x = 6 x^2 = 6*6 = 36 a = 144/36 = 4 which means equation is 4x^2 therefore, the answer is c

random231 (random231):

 do you get till here.